Evaluate the integral of √x dx/1+x√x


1
Expert's answer
2021-02-24T12:38:52-0500

"\int\frac{\sqrt xdx}{1+x\sqrt x}"

Solution:

"I=\int\frac{\sqrt xdx}{1+x\sqrt x}"

"\sqrt x=t;" "x=t^2;" "dx=2tdt" .

"I=\int \frac{2t^2dt}{1+t^3}"

"\frac{2t^2dt}{1+t^3}=\frac{2t^2}{(t+1)(t^2-t+1)}=\frac{2t^2}{(t+1)((t-\frac12)^2+\frac34)}"

Expand the fraction:

"\frac{2t^2}{(t+1)((t-\frac12)^2+\frac34)}=\frac{2}{3(t+1)}+\frac{\frac43(t-\frac12)}{(t-\frac12)^2+\frac34}"

"I=\int (\frac{2}{3(t+1)}+\frac{\frac43(t-\frac12)}{(t-\frac12)^2+\frac34})dt="

"\int \frac{2dt}{3(t+1)}+\int\frac{\frac23d(t-\frac12)^2}{(t-\frac12)^2+\frac34}="

"\frac23(\ln{|t+1|}+\ln{{((t-\frac12)^2+\frac34})})+C="

"\frac23\ln{|1+t^3|}+C=\frac23\ln{|1+x\sqrt x|}+C=\ln{(1+x\sqrt x)^{\frac23}}+C"

Answer: "\int\frac{\sqrt xdx}{1+x\sqrt x}=\ln{(1+x\sqrt x)^{\frac23}}+C"